# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ 1.2 Mountain pine beetle life history traits
A key element of this paper is a nonlinear dynamical model that captures key MPB life history traits. In this section, we review the traits with a view towards qualitative features that should be emphasized in the model. Here, life history traits refers to the nonstandard aspects of MPB’s host-parasitoid interaction with lodgepole pine, including the ways in which MPB “decides” about what trees to attack, and how MPB responds to variation in the density of host trees and conspecific beetles. Our list of life history traits is not exhaustive. Previous research has identified some traits that do not show up as a clear pattern in our data, or traits that we cannot feasibly incorporate into our models. These are discussed further in Appendix B.
It is worth keeping in mind that what counts as a life history trait is somewhat a matter of perspective. As we will discuss, both MPB’s stand-dependent dispersal and large-tree attack preference could potentially stem from its inability to chemically sense small trees; are these two separate life-history traits, or one combined trait? We take a pragmatic approach, favoring traits that are conceptually separable and traits that can be independently manipulated in our mechanistic model.
@@ 1.2.1 Allee effect
Mountain pine beetle (MPB) is subject to an Allee effect, a phenomenon where individual fitness increases with population density. The Allee effect arises because a threshold number of beetles must attack a single tree to successfully kill it, but this threshold is challenging to reach in low-density populations. Another common mechanism of an Allee effect, poor mate finding in small populations, has been ruled out by previous research (Amman,, 1980; Bleiker et al.,, 2014).
The Allee effect is mitigated by MPB’s propensity for aggregation. In general, aggregation counteracts the negative effects of positive density-dependence, just as dispersal counteracts the negative effects of negative density dependence (Crowley,, 1981; Briggs and Hoopes,, 2004). Females release aggregation pheromones and attacked trees emit terpene volatiles, which together attract additional beetles. However, this chemical signal is dampened by unsuitable host-tree chemistry (Hynum and Berryman,, 1980; Boone et al.,, 2011), unsuitable natal-tree chemistry (which negatively affects aggregation pheromone production; Chiu et al.,, 2018), and a limited detection limit for aggregation pheromones, which is around 100-1000 meters (Barclay et al.,, 1998, Fig. 2; Strohm et al.,, 2013, Fig. 4; Mitchell and Preisler,, 1991, Fig. 4–5). Thus, the Allee effect persists due to various physical and chemical factors that constrain the strength of aggregation.
@@
There are at least three explanations for why the Allee effect is not evident in some datasets. 1) Within-tree beetle density is positively related to the probability of tree death and the likelihood of beetle reproduction, but the counter-veiling impact of negative density dependence cancels out this signal. Beetle density at the tree level is negatively related to per capita oviposition and per capita survival (Raffa and Berryman,, 1983, Fig. 10). Further, at the stand level, high beetle density suggests that many trees have been killed and that predominantly low-quality trees remain. Thus, the demographic Allee effect at the stand level is weaker than the component Allee effect at the tree level. 2) While aggregation counteracts the negative effects of positive density-dependence, it also makes it difficult to observe a low-density but non-zero population at the stand level. To see this, consider the hypothetical scenario where aggregation was perfect (i.e., beetles can travel and communicate perfectly over long distances). In this scenario, you would never observe a tree with a low attack density, so you could never uncover a positive relationship between attack density and tree death. Indeed, tree-level studies fail to uncover an Allee effect when they only examine successfully attacked trees (Reid,, 1963; Berryman,, 1974, Fig. 2c; Berryman,, 1976, Fig. 2). We don’t live in the perfect aggregation scenario — in reality, we observe some trees with a sub-threshold attack density — but perhaps aggregation is strong enough that we rarely observe stands in which the overall beetle density is below the theoretical stand-level allee threshold. 3) The majority of beetles are thought to disperse less than 30 meters, so by calculating the density of infested trees over acres or hectares, the aforementioned studies (MacQuarrie and Cooke,, 2011; Tunnock et al.,, 1970; Trzcinski and Reid,, 2009) are calculating a spatial average over a number of weakly coupled sub-populations. Averaging effectively reduces the variation in beetle density, and in turn, makes it more difficult to detect an Allee effect. This statistical stabilization hypothesis has been used to explain weak population regulation across many species (De Roos et al.,, 1991; Ray and Hastings,, 1996; Thorson et al.,, 2015).
1.2.2 Stand-Dependent Dispersal
@@
MPB predominantly disperses from stands only after most large-diameter trees are killed. The vast majority of these emigrants die on their journey, either due to an Allee effect or a low energetic condition at their destination. Since stand-dependent dispersal and the associated mortality are somewhat obscure features of MPB dynamics, we will take a moment to discuss supporting evidence.
There are several potential explanations for stand-dependent dispersal. One hypothesis suggests that MPB is simply unable to aggregate effectively on small and medium-sized trees, putatively due to chemosensory constraints. Field data support this — MPB primarily attacks small trees when they are adjacent to large trees (Mitchell and Preisler,, 1991; Preisler and Mitchell,, 1993), analogously to how MPB attacks nonhost trees only when they are adjacent to heavily-infested host trees (Amman et al.,, 1985; Huber et al.,, 2009). Further, studies have shown large trees have higher monoterpene concentrations, which may be the main determinant of whether pioneer female beetles initiate an attack (Boone et al.,, 2011; Seybold et al.,, 2006). A secondary explanation considers how beetle infestations alter the forest microclimate (Bartos,, 1989; Amman and Logan,, 1998). As stands are thinned, increased wind disrupts MPB aggregation pheromone plumes, and increased light may induce MPB to fly. A final hypothesis claims that beetle stand-dependent dispersal is a behavioral adaptation that maximizes inclusive fitness. Large trees have thicker phloem leading to high per capita fecundity, so lineages that seek out large trees may have higher geometric fitness than their less-risky competitors, despite the high probability of dispersal mortality. To our knowledge, this explanation has not yet been explored by the literature.
We propose that stand-dependent dispersal in beetles is best explained by the chemical-sensing hypothesis, i.e., beetles are less effective at mass attacking small trees. This tentative conclusion was reached by a process of elimination. The problem with the microclimate hypothesis is that there is a time lag between a beetle infestation and biotic forest thinning. Most MPB-killed trees don’t fall down for 5-20 years (Lewis and Hartley,, 2006; Hansen,, 2014), and most don’t lose their needles for 2–7 years (Klutsch et al.,, 2009; Audley et al.,, 2021), but the populations crashes occur 1–2 years after infestations have peaked (Fig. D.3). The inclusive fitness hypothesis is unlikely because disperses face significant mortality, and because beetles have positive expected fitness when they do successfully kill medium-sized trees (Fig. 4).
@@
Even though we cannot track individual beetle dispersal, there is robust evidence for stand-dependent dispersal. In addition to the previously discussed evidence, pheromone-trap experiments show substantially higher numbers of trapped beetles in thinned vs. unthinned stands (Bartos,, 1989; Schmitz et al.,, 1989). Sallé and Raffa, (2007) have shown that beetles with low body weight had proportionally more lipids. Since beetles are more reactive to aggregation pheromones once lipids are depleted, this explicates a physiological mechanism by which beetles can escape depleted stands. Several modelling studies explicitly infer substantial stand-dependent dispersal(Powell and Bentz,, 2014; Powell et al.,, 2018; Howe et al.,, 2021). Other, more complicated models implicitly assume stand-dependent dispersal, and then show that the model generates realistic behavior (e.g., Powell et al.,, 2018; Bone and Altaweel,, 2014; Kautz et al.,, 2016).
@@
The most compelling evidence for stand-dependent dispersal is the fact that outbreaks invariably end before the vast majority of medium-sized trees are killed (Fig. 5), even though beetles have positive fitness on medium-sized trees (Fig. 4). Previous research (Goodsman et al.,, 2018) has explained this pattern —– where an outbreak ends sooner than expected —– as a result of overcompensatory density dependence, a phenomenon where increased population size now leads to a decreased population size in the next generation. However, there is not strong evidence for overcompensatory density dependence in MPB. Attacking beetles pack themselves into trees at a density that optimizes individual fitness (Raffa and Berryman,, 1983). When attack densities do exceed the theoretical optimum, there is no clear change in emergence densities (Fig. D.2; Berryman,, 1974, Fig. 2c; Berryman,, 1976, Fig. 2, Raffa and Berryman,, 1983, Fig. 6 shows a slight increase) or population dynamics (Appendix A.4).
@@
Figure 4: Mountain pine beetle fitness as a function of diameter at breast height (DBH), in inches. Ecological fitness is estimated with (0.3)×(0.3$\times$( 0.3 ) × females out of the tree / females into the tree, where the factor $0.3$ is the approximately the median dispersal-phase survival in the dataset from Klein et al., (See Fig. D.1). Depending on the study, gallery starts or entry holes serve as estimates for females into the tree (note that only females initiate galleries). (2/3)×(3/2)×(2/3$\times$(3/2$\times$( 2 / 3 ) × ( 3 / 2 ) × Emergence holes, or $(2/3)$ of the overwintered brood serve as estimates of females out of the tree. Note that (2/3) is the typical proportion of females (Reid,, 1962; Cole et al.,, 1976), and that (3/2) is the average number of beetles emerging from each hole (Safranyik and Linton,, 1985; Peterman,, 1974, Fig. 13). Data comes from Goodsman et al., (2018); Cudmore et al., (2009); Reid, (1963); Safranyik and Carroll, (2006). The means and standard errors were only computed for bins with at least 5 observations.
@@
Figure 5: The proportion of remaining live trees as a function of time, stratified by diameter class (diameter at breast height, DBH). Data comes from Stipe, (1974, 1976); Gibson and Bennett, (1978); Gibson et al., (1979); Klein et al., (1978); Preisler and Mitchell, (1993); Nelson, (2009).
@@
There is also evidence that stand-dependent dispersal mostly results in mortality, although this evidence is indirect. First, it is our observation (based on Aerial Overview Survey data) that outbreaks do not last substantially longer in areas with high contiguous pine biomass. If dispersing beetles were not dying in such areas, emigrating beetles would be replaced by immigrating beetles, and thus outbreaks would be prolonged. Second, mortality is already substantial for beetles traveling short distances: baseline dispersal mortality is estimated to be around 40–80% (Safranyik,, 1989; Latty,, 2007; Pope et al.,, 1980); Fig. D.1). Third, dispersal puts beetles in a low energetic condition, leading to decreased aggregation pheromone production, and making them less likely to attack trees (Jones,, 2019; Jones et al.,, 2020). Fourth, lab experiments indicate that most beetles only survive several days outside of a tree (Fuchs and Borden,, 1985; Reid and Reid,, 2008).
1.2.3 Size-dependent attack preference
@@
Large trees offer several benefits to MPB, so it is no surprise that MPB has a strong preference for large-diameter trees in the landing, pioneer attack, and mass attack phases of its life cycle. In experiments, beetles prefer landing on large, vertically-oriented black silhouettes (Shepherd,, 1966; Billings et al.,, 1976), indicating that visual cues play a major role in initial, pre-aggregation landings. Chemo-sensory cues play a secondary role in the initial landing phase (Billings et al.,, 1976; Hynum and Berryman,, 1980; Moeck and Simmons,, 1991). After landing, a female pioneer beetle decides whether to attack based on a combination of physical and gustatory cues associated with tree diameter, such as terpene concentrations (Raffa and Berryman,, 1982) and bark roughness (Reid,, 1963). Once mass attacks have been initiated, both males and females are attracted to large trees, due to a synergist attraction to aggregation pheromones and volatilized terpenes, both of which scale with tree size.
@@
MPB’s large-tree attack preference is extremely well-established. The probability of tree death increases logistically with tree diameter (Buonanduci et al.,, 2023; Rodman et al.,, 2022; Morris et al.,, 2022; Negrón,, 2020; Negrón et al.,, 2017; Björklund and Lindgren,, 2009 and 15 sources therein, see Table 1), and stand susceptibility indexes invariably use variables that correlate with tree diameter: tree age, pine volume, or basal area (Shore et al.,, 1992; Chojnacky et al.,, 2000; Randall et al.,, 2011; Bentz et al.,, 1993, and sources therein). Depending on site-specific factors, the probability of tree mortality is near-zero for 4–8-inch trees and near-100% for 12–16-inch trees.
1.2.4 Size-dependent fecundity
@@
MPB has higher effective per capita fecundity in larger trees. Large trees have thick phloem, which is the primary food source of larvae and teneral adults. Thin phloem can slow larval development, which can cause mortality if a cold snap occurs before the cold-resistant third & fourth-instar stages are attained; if larvae do manage to fully develop within a small tree, they may have small body sizes and low energetic condition upon emergence (Amman,, 1983). Additionally, the thick bark of large trees protects against cold winter temperatures.

### v2

@@ 1.2 Mountain pine beetle life history
A key element of this paper is a nonlinear dynamical model that captures key features of MPB biology. In this section, we review life history with a view toward features that should be emphasized in the model, like the ways in which MPB “decides” about what trees to attack, and how MPB responds to variation in the density of host trees and conspecific beetles. Our list is not exhaustive. Previous research has identified some biological processes that do not show up as a clear pattern in our data, or that we cannot feasibly incorporate into our models. These are discussed further in Appendix B.
@@ 1.2.1 Allee effect
Mountain pine beetle is subject to a strong Allee effect, a phenomenon where per capita growth rates become negative at a sufficiently low population density. The Allee effect arises because a threshold number of beetles must attack a single tree to successfully kill it. Another common mechanism of an Allee effect, poor mate finding in small populations, has been ruled out by previous research (Amman,, 1980; Bleiker et al.,, 2014).
The Allee effect is mitigated by MPB’s ability to aggregate. In general, aggregation counteracts the negative effects of positive density dependence, just as dispersal counteracts the negative effects of negative density dependence (Crowley,, 1981; Briggs and Hoopes,, 2004). Aggregation pheromones and volatile monoterpene from attacked trees attract additional beetles. However, this chemical signal may be dampened by unsuitable host-tree chemistry (Hynum and Berryman,, 1980; Boone et al.,, 2011), unsuitable natal-tree chemistry (which negatively affects aggregation pheromone production; Chiu et al.,, 2018), or simply a limited ability to detect aggregation pheromones at-distance. Evidence from small-scale spatial maps, as well as mechanistic pheromone diffusion models, suggest that beetles primarily respond to aggregation pheromones that are 20-100 meters from their tree-of-origin ( Mitchell and Preisler,, 1991, Fig. 4–5); Strohm et al.,, 2013). There is limited evidence that the detection limit is 500-1000 meters under the right wind conditions (Barclay et al.,, 1998). In summary, the Allee effect persists due to various physical and chemical factors that constrain the strength of aggregation.
@@
1.2.2 Stand-Dependent dispersal mortality
@@
Mountain pine beetle predominantly disperses from stands only after most large-diameter trees are killed. The vast majority of these emigrants die on their journey due abiotic elements or an inability to find a suitable host tree. Beetles that do survive may be less likely to attack suitable host trees, given the low energetic conditions resulting from a long flight (Jones et al.,, 2020). Since stand-dependent dispersal and the associated mortality are somewhat obscure features of MPB dynamics, we will take a moment to discuss supporting evidence.
The most likely explanataion for stand-dependent dispersal is chemosensory constraints: MPB is simply unable to sense and aggregate on small and medium-sized trees. Multiple lines of evidence support this explanation — MPB primarily attacks small or non-host trees when they are adjacent to large host trees (Mitchell and Preisler,, 1991; Preisler and Mitchell,, 1993; Amman et al.,, 1985; Huber et al.,, 2009), and studies have shown large trees have higher monoterpene concentrations, which may be the main determinant of whether pioneer female beetles initiate an attack (Boone et al.,, 2011; Seybold et al.,, 2006; Howe et al., 2022a, ).
Alternative explanations for stand-dependent dispersal have less support. One considers the fact that beetle infestations alter the forest microclimate (Bartos,, 1989; Amman and Logan,, 1998). As stands are thinned, increased wind disrupts MPB aggregation pheromone plumes, and increased light may induce MPB to fly. However, the microclimate hypothesis is undermined by timing — trees take 5-20 years to fall (Lewis and Hartley,, 2006; Hansen,, 2014) and 2–7 years to lose needles (Klutsch et al.,, 2009; Audley et al.,, 2021), while populations crash within 1–2 years (Fig.,E.1). The inclusive fitness hypothesis suggests beetles refuse to consider medium-sized trees because large trees engender higher fitness, but this is unlikely given high dispersal mortality and positive fitness in medium trees (Fig.,3).
@@
Even though we cannot track individual beetle dispersal, there is robust evidence for stand-dependent dispersal. In addition to the previously discussed evidence, trapping experiments have shown that far more beetles are caught in unthinned stands than thinned stands, regardless of whether pheromone traps (Bartos,, 1989; Schmitz et al.,, 1989), or passive traps are used (Fig. 2). Sallé and Raffa, (2007) have shown that beetles with low body weight had proportionally more lipids. Since beetles are more reactive to aggregation pheromones once lipids are depleted, this explicates a physiological mechanism by which beetles can escape depleted stands. Several modeling studies explicitly infer substantial stand-dependent dispersal(Powell and Bentz,, 2014; Powell et al.,, 2018; Howe et al.,, 2021). Other, more complicated models implicitly assume stand-dependent dispersal, and then show that this assumption leads to realistic population dynamics (e.g., Powell et al.,, 2018; Bone and Altaweel,, 2014; Kautz et al.,, 2016).
Figure 2: Fewer beetles are trapped in heavily thinned stands, suggesting that beetles disperse away from sparse forests. The y-axis displays the mean and standard errors of beetles in passive traps in Fraser Experimental Forest’s (Colorado) managed stands over the full flight period. Thinning levels are categorized as ’Low’, ’Medium’, and ’High’, with corresponding Growth Stocking Levels (GSL) of 120, 80, and 40, respectively. The small size of the experimental blocks (0.2 ha) suggests that captured beetles predominantly originated from adjacent unmanaged forest, in turn suggesting that the data reflect the impact of tree density on dispersal behavior, not on the within-block infestation density. Data comes from Negrón, (2018).
@@
The most compelling evidence for stand-dependent dispersal is the fact that outbreaks invariably end before the vast majority of medium-sized trees are killed (Fig. 4), even though beetles have positive fitness on medium-sized trees (Fig. 3). Previous research (Goodsman et al.,, 2018) has explained this pattern —– where an outbreak ends sooner than expected —– as a result of overcompensatory density dependence, a phenomenon where increased population size now leads to a decreased population size in the next generation. However, there is no evidence for overcompensatory density dependence in MPB. Attacking beetles pack themselves into trees at a density that roughly optimizes individual fitness (Raffa and Berryman,, 1983). When attack densities do exceed the theoretical optimum, there is no clear change in emergence densities (Fig. E.3; Berryman,, 1974, Fig. 2c; Berryman,, 1976; Fig. 2, Raffa and Berryman,, 1983, Fig. 6 shows a slight increase) and there is no clear change in population dynamics (Appendix A.6).
@@
Figure 3: Mountain pine beetle tends to have positive fitness when attacking medium-sized trees (approx. 8–14 inches DBH). Ecological fitness is estimated with (0.3)×(0.3$\times$( 0.3 ) × females out of the tree / females into the tree, where the factor $0.3$ is the approximately the median dispersal-phase survival in the dataset from Klein et al., (See Fig. E.4). Depending on the study, gallery starts or entry holes serve as estimates for females into the tree (note that only females initiate galleries). (2/3)×(3/2)×(2/3$\times$(3/2$\times$( 2 / 3 ) × ( 3 / 2 ) × Emergence holes, or $(2/3)$ of the overwintered brood serve as estimates of females out of the tree. Note that (2/3) is the typical proportion of females (Reid,, 1962; Cole et al.,, 1976), and that (3/2) is the average number of beetles emerging from each hole (Safranyik and Linton,, 1985; Peterman,, 1974, Fig. 13). Data comes from Goodsman et al., (2018); Cudmore et al., (2009); Reid, (1963); Safranyik and Carroll, (2006). The means and standard errors were only computed for bins with at least 5 observations.
@@
Figure 4: Outbreaks typically end before a majority of medium-sized trees (approx. 8–14 inches DBH) are killed. The figure depicts the proportion of remaining live trees as a function of time, stratified by diameter class (diameter at breast height, DBH). Data comes from Stipe, (1974, 1976); Gibson and Bennett, (1978); Gibson et al., (1979); Klein et al., (1978); Preisler and Mitchell, (1993); Nelson, (2009).
@@
There is also evidence that stand-dependent dispersal mostly results in mortality, although this evidence is indirect. First, it is our observation (based on Aerial Overview Survey data) that outbreaks do not last substantially longer in areas with high contiguous pine biomass. If dispersing beetles were not dying in such areas, emigrating beetles would be replaced by immigrating beetles, and thus outbreaks would be prolonged. Second, mortality is already substantial for beetles traveling short distances: baseline dispersal mortality is estimated to be around 40–80% (Safranyik,, 1989; Latty,, 2007; Pope et al.,, 1980); Fig. E.4). Third, dispersal puts beetles in a low energetic condition, leading to decreased aggregation pheromone production, and making them less likely to attack trees (Jones,, 2019; Jones et al.,, 2020). Fourth, lab experiments indicate that most beetles only survive several days outside of a tree (Fuchs and Borden,, 1985; Reid and Reid,, 2008). Fifth, longer flights increase exposure to natural enemies, most notably avian predators (Rust,, 1930; Stallcup et al.,, 1963).
1.2.3 Large-tree preference
Mountain pine beetle exhibits a strong preference for large-diameter trees during their dispersal/host-selection phase, driven by both proximate sensory cues and ultimate evolutionary benefits. From an evolutionary perspective, large trees offer two key advantages: primarily, they have thicker phloem, which serves as the primary food source for MPB larvae, and secondarily, their thicker bark may provide better insulation during cold winters.
@@ 1.2.3 Large-tree preference
The proximate mechanisms driving a large-tree preference are evident across different stages of host selection. In the initial landing phase, beetles display a mix of random movement and slight preference for large, vertically-oriented black silhouettes (Shepherd,, 1966; Billings et al.,, 1976), indicating that visual cues play a role. Chemo-sensory cues play a relatively minor role during the initial landing phase (Billings et al.,, 1976; Hynum and Berryman,, 1980; Moeck and Simmons,, 1991). For pioneer attacks, females use physical and gustatory cues associated with tree diameter, such as terpene concentrations (Raffa and Berryman,, 1982) and bark roughness (Reid,, 1963). During mass attacks, larger trees become even more attractive due to increased concentrations of aggregation pheromones and host volatiles, both of which correlate positively with tree size. These sensory preferences align with the evolutionary advantages of large trees, ensuring that beetles select hosts that offer the best resources for their offspring.
@@
Mountain pine beetle’s large-tree attack preference is extremely well-established. The probability of tree death increases logistically with tree diameter (Buonanduci et al.,, 2023; Rodman et al.,, 2022; Morris et al.,, 2022; Negrón,, 2020; Negrón et al.,, 2017; Björklund and Lindgren,, 2009 and 15 sources therein, see Table 1), and stand susceptibility indexes invariably use variables that correlate with tree diameter: tree age, pine volume, or basal area (Shore et al.,, 1992; Chojnacky et al.,, 2000; Randall et al.,, 2011; Bentz et al.,, 1993, and sources therein). Depending on site-specific factors, the probability of tree mortality is near-zero for 4–8-inch trees and near-100% for 12–16-inch trees.
Despite their preference for large trees, MPB have developed strategies to maintain a near-optimal attack density and avoid strong intraspecific competition. Mountain pine beetle will “fill up” trees to an approximately invariant density, measured in beetles per square meter of tree trunk surface area, before moving on to the next tree. Interestingly, this attack density is close to being optimal in terms of the population growth rate (Raffa and Berryman,, 1983). The mechanism for attaining (approximate) optimality is a suite of aggregation and anti-aggregation pheromones, and the relationship between such pheromones and host tree defenses. Beetles are attracted to the combination of terpenoid aggregation pheromones and volatilized terpenes. Because both are derived from tree resin, trees become less attractive as their resin defenses are depleted (Raffa and Berryman,, 1983; Raffa,, 1988). Anti-aggregation pheromones, which are only detected by beetles at short distances (around a meter; Bentz,, 1996), redirect beetles to nearby trees, while rivalry stridulations (noises produced by males) help ensure that galleries are evenly spaced. Raffa, (2001) puts forward several hypotheses for the apparent lack of cheaters — beetles who ignore the anti-aggregation pheromones and enter a “full” tree.
Beetle attack density remains relatively constant across tree diameters and beetle population sizes. Killed trees receive an approximately invariant number of attacks per unit surface area (Fig.,A.9, Reid,, 1963, Table 1, Peterman,, 1974, Figs. 4–7). There is less literature on the relationship between attack density and beetle population size, but any such relationship is likely weak, given that 1) Negrón, (2018) found no relationship; 2) there is a clear mechanism for limiting attack density (see the previous paragraph), and 3) that attack density varies in the range of 50–100 females per square meter (see Fig. E.5 in the Results), despite beetle population sizes varying by several orders of magnitude.
1.2.4 Tree size-dependent fecundity
@@
Mountain pine beetle has higher effective per capita fecundity in larger trees. Large trees have thick phloem, which is the primary food source of larvae and teneral adults. Thin phloem can slow larval development, which can cause mortality if a cold snap occurs before the cold-resistant third & fourth-instar stages are attained; if larvae do manage to fully develop within a small tree, they may have small body sizes and low energetic condition upon emergence (Amman,, 1983). Additionally, the thick bark of large trees protects against cold winter temperatures.